y-(-5)=2/3(x-7)
the answer is
y=2/3x-29/3
PlEASE tell me how they get this answer where does the
-29 come from??
tthank you

Answer by Alan3354(69443)

y-(-5)=2/3(x-7)
-----------
y + 5 = (2x-14)/3
5 = 15/3
y = 2x/3 - 14/3 - 15/3
y = 2x/3 - 29/3